Fermi National Laboratory
 

Accelerator Update for June 28, 2006 - June 29, 2006

 

Wednesday June 28

The day shift began with Operations monitoring store 4802 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with MiniBooNE and NuMI taking beam.

Recycler experts turned off the Pelletron at 8:16 AM so that air conditioning technicians could perform maintenance.

Operations prepared MiniBooNE for access at 8:43 AM. Experts replaced a fuse for the dehumidifier system. MiniBooNE resumed taking beam at 9:52 AM.

The Recycler air conditioning work was completed around 12 PM. The Pelletron was turned back on.

Operations began preparing the Recycler at 12:24 PM to stash antiprotons. Operations reported at 1:10 PM that the Recycler had taken one transfer, but then there was a problem with the low level RF system.

At 1:30 PM, Operations reported that the P1 transfer line was down due to a Lambertson power supply problem (LAM52). This also held off beam to NuMI. Operations called an expert.

The stash was canceled and Pbar resumed stacking at 2:44 PM. NuMI also resumed taking beam and the Meson MTest experiment began taking beam.

At 2:45 PM, Operations reported that Booster was not cogging correctly. Experts were tuning and working with the RF at the time.

At 6:52 PM, a Switchyard (SY) power supply (HP3DS) tripped off and pulled the beam permit. This was only the first of many times this supply tripped during the evening. Operations called in an expert to investigate.

The Recycler prepared to stash at 7:34 PM. The stash ended and Pbar resumed stacking at 8:47 PM.

Operations terminated store 4802 at 9:08 PM.

Operations prepared D0 for an access at 9:23 PM.

At 10:25 PM, the H- Source tripped off. It reset okay.

Operations put the TeV into shot setup at 10:45 PM. This shut off the MTest beam.

Thursday June 29

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with normal Linac, MI, Recycler, and Booster operations, and with MiniBooNE and NuMI taking beam.

Operations established store 4804 at 1:12 AM with an initial luminosity of 117.67E30. Pbar resumed stacking.

At 1:38 AM, the quadrupole magnet power supply for a Linac RF station (LRF3) and a vacuum pump (LIP600) tripped off and then the vacuum pump tripped off two more times over the next thirty minutes. Both device reset okay.

At 1:43 AM, the NuMI Horn began tripping off. A NuMI expert told Operations to reset everything. He believed the problem was related to flaking nickel.

The Recycler prepared to stash at 1:58 AM. The stash ended. Pbar resumed stacking and NuMI resumed taking beam at 2:10 AM.

At 6:08 AM, Operations reported that the CDF luminosity had dropped to zero. Experimenters reported that their solenoid had tripped off due a low liquid helium level.

The Recycler prepared to stash at 7:20 AM.

The Plans for Thursday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/29/2006    email Fermilab